# Fluid Antenna-enabled Near-Field Integrated Sensing, Computing and Semantic Communication for Emerging Applications

### v1

@@ Abstract
Index Terms:
Integrated sensing and communication, semantic communication, transmit beamforming, and fluid antennas.
@@ I Introduction
With the rapid growth of emerging applications such as virtual/augmented reality (VR/AR) and digital twins (DTs), the integration of wireless sensing and communication is ready to play a pivotal role in enabling next-generation services. Recognising its transformative potential, the International Telecommunication Union (ITU) has incorporated integrated sensing and communication (ISAC) as one of the six core usage scenarios in its global vision for the sixth-generation (6G) mobile communication systems.
@@
The basic principle of ISAC is to deploy wireless sensing and communication capability through the shared use of wireless resources, including power, frequency bands, beams, and hardware infrastructure. Numerous studies, to list a few, [1, 2, 3, 4], have demonstrated that ISAC systems can outperform standalone sensing and communication systems in terms of resource efficiency and can unlock synergistic advantages, where sensing enhances communication and communication supports sensing. On the one hand, sensing-assisted communication uses sensing data to provide contextual awareness, which helps optimise communication processes like beamforming, channel estimation, and resource allocation by inferring channel quality and proactively modifying the system’s settings [5]. On the other hand, in a communication-assisted sensing framework, communication devices support sensing by transmitting sensory data and capturing reflected signals, eliminating the need for dedicated sensors. This simplifies network architecture and reduces hardware costs compared to traditional sensor-based networks [6].
[…]
@@
Several key advantages emerge when considering the NF model. First, unlike far-field (FF) beamforming, which directs beam energy along a fixed path, NF beam focusing utilises spherical wavefronts to concentrate energy at a specific location [9, 10]. This focusing capability enhances the received signal power for intended users. In terms of sensing capabilities, the NF model demonstrates superior performance in target localisation [11]. The spherical wavefront properties enable simultaneous estimation of both angular and range information with high precision, particularly when employing advanced techniques such as multiple signal classification (MUSIC) with fine-grid processing. While Doppler-based methods in FF scenarios can estimate range through velocity measurements, those methods often yield less accurate results, especially in static or low-mobility scenarios where frequency shifts are minimal. Furthermore, the NF beam focusing effect provides enhanced control over the sensing signal distribution, effectively reducing interference in echo signals and improving overall sensing accuracy [12]. This characteristic is particularly valuable in complex environments where multiple line-of-sight (LoS) targets or reflectors may be present. Several studies have explored the NF-ISAC system. For instance, Wang et al. [13] analysed NF-ISAC systems with multiple communication users and a single target, investigating both fully digital antenna arrays and hybrid digital-analogue configurations. Their findings highlight the performance enhancements enabled by exploiting the additional distance dimension in NF-ISAC compared to FF-ISAC. Zhao et al. [14] examined the trade-off between communication and sensing functionalities in both downlink and uplink scenarios. Qu et al. [15] compared NF beam focusing with FF beam steering, demonstrating through simulations that NF beam focusing effectively mitigates co-angle interference while enhancing both communication and sensing performance.
@@
While NF-ISAC offers promising advantages, it faces substantial challenges, particularly in managing the enormous volume of data transmitted and received in real-time scenarios such as autonomous driving. The excessive data volume can lead to increased latency and computational overhead, hindering system performance [16]. Another critical challenge involves potential data privacy risks. As sensing and communication signals are transmitted simultaneously, sensing targets (which are unintended communication users) may inadvertently receive confidential communication messages, raising significant privacy concerns [17]. The integration of semantic communication into the NF-ISAC framework presents a promising solution to these challenges. Semantic communication focuses on transmitting the meaning and relevance of information rather than raw data, enabling intelligent prioritisation and compression based on contextual importance [18]. For example, in autonomous driving, instead of transmitting comprehensive sensing results, the system could extract and transmit semantically relevant information, such as the trajectories of critical objects or potential hazards, to other roadside units. This semantic-aware approach significantly reduces the data burden and enhances system efficiency by ensuring that limited resources are allocated to transmitting the most meaningful and actionable information. To enable semantic communication, both transmitters and receivers must establish shared knowledge bases (KBs) containing essential information accessible to both parties [19, 20]. Crucially, the security and integrity of these KBs are central to maintaining the privacy and security of semantic communication. Regulations such as the General Data Protection Regulation (GDPR), which governs the handling of personal data within the EU and EEA, provide a legal and procedural framework for protecting sensitive data [21, 22]. By adhering to GDPR principles or similar regulatory frameworks, KBs can be safeguarded against unauthorised access and cyber-attacks [23, 24]. As a result, semantic communication inherently enhances data privacy: since only high-level semantic representations are transmitted, and successful decoding requires access to valid KBs, unintended users are effectively excluded from recovering the original message. Moreover, semantic communication facilitates efficient multi-modal data transmission, meeting the diverse requirements of different communication users [25]. Despite these advantages of semantic communication, existing research on semantic communication, such as [26, 27, 28], has largely overlooked its integration into the NF-ISAC framework. This research gap represents a significant opportunity for advancing NF-ISAC systems by leveraging the benefits of semantic communication to address data overload, privacy concerns, and multi-modal data requirements.
@@
Nonetheless, integrating semantic communication into the NF-ISAC framework introduces specific challenges, particularly due to the interference caused by the sensing signals. Unlike classical communication, which focuses on bit-level accuracy, semantic communication aims to preserve the intended meaning or context of the transmitted information. This makes it inherently more sensitive to channel conditions and interferences. Under poor channel conditions, not only can the data become corrupted, but the semantic meaning of a message may also be distorted—a phenomenon referred to as semantic noise [18]. Semantic noise occurs when the receiver misinterprets the message, deviating from the sender’s original intent, which can lead to misunderstandings or incorrect decisions. For example, suppose a sender intends to transmit the message “Alice’s new bike was broken”. In a traditional communication system, a noisy channel might corrupt it to “Alace’s new baki was brakon”, maintaining the structure but affecting intelligibility. In contrast, a semantic communication system might deliver “Alisa’s bike is broken,” preserving a plausible meaning but deviating from the exact intended context. Therefore, developing strategies to mitigate semantic noise is essential for maintaining reliability in semantic communication systems.
@@
One promising technique to stabilise the channel condition is to use the fluid antennas (FAs), which are flexible in terms of their positions to reconfigure the radiation characteristics [29, 30, 31]. Compared to fixed-position antennas (FPAs), the FAs have a greater degree of freedom (DoF) to explore the channel variations to adapt to the changing channel conditions. Various studies have highlighted the significant advantages of FAs in wireless communication systems. For instance, Zhu et al. [32] compared the communication performance of FPAs and FAs by analysing the maximum channel gain achieved in deterministic and stochastic channel environments. Their findings show that FAs outperform FPAs, especially in scenarios with an increasing number of channel paths, due to their ability to exploit more pronounced small-scale fading effects in the spatial domain. Similarly, Wang et al. [33] explored the role of FAs in ISAC systems within a multi-user multiple-input multiple-output (MU-MIMO) downlink framework. In this study, a base station (BS) equipped with FAs demonstrated enhanced beamforming precision, reduced inter-user interference, and improved overall system capacity. These advancements illustrate the potential of FAs to optimise both communication and sensing performance in complex multi-user scenarios. Despite the substantial progress made in demonstrating the benefits of FAs for improving communication performance in conventional systems, their applications in semantic communication and NF-ISAC systems have received minimal attention.
@@
In [34], the framework of integrated sensing, computing, and semantic communication (ISCSC) was proposed. However, the impact of NF effects and the potential advantages of incorporating FAs into the system were not explored. This limitation highlights a significant research gap, motivating the development of fluid antenna-enabled NF-ISCSC (FA-enabled NF-ISCSC) systems. Compared to the ISAC design, the proposed framework achieves higher data rates without sacrificing sensing performance, or alternatively, enhances sensing accuracy while maintaining the same data rate. These advantages make it especially suitable for applications such as vehicular networks and digital twins. In vehicular networks, particularly within the intelligent transportation system (ITS), sensing enables roadside units (RSUs) and vehicles to acquire essential environmental information, such as traffic signs, road conditions, and dynamic obstacles [35]. Semantic communication ensures that only task-relevant and context-specific information, such as the sudden appearance of a pedestrian, is exchanged among RSUs and between RSUs and vehicles. For example, rather than transmitting raw sensor data, an RSU can convey a high-level semantic message like “vehicle merging from left at 25 m/s”, thereby reducing the communication load and improving the accuracy and timeliness of information delivery. The use of FAs further enhances communication reliability and adaptability by reducing semantic errors. Computing plays a critical role by processing sensed data for real-time scene understanding, including object detection, tracking, and trajectory prediction [36]. Additionally, computing enables semantic communication by extracting and prioritising key features from sensory data for efficient and meaningful transmission. In the context of digital twins, sensing data collected by end devices enables real-time perception of the physical entity’s state. In the uplink scenario, semantic communication allows the end device to interpret and filter out irrelevant information prior to the transmission, thereby reducing bandwidth consumption and minimising latency when uploading to the digital twin server [37]. In the downlink, the server can generate and transmit semantic information based on environmental context, thus alleviating downlink resource congestion. For example, in the context of a digital twin for eHealth, instead of transmitting the full message “The patient is a 68-year-old male with a medical history of Type 2 diabetes mellitus, hypertension, and coronary artery disease. He presented to the emergency department with chest pain that radiated to his left arm”, the end device may send a high-level semantic interpretation such as “68-year-old male with chronic cardiovascular and metabolic diseases presented with radiating chest pain to the ED”. In this semantically extracted message, the word “chronic” captures the notion of a long-standing medical history, Type 2 diabetes mellitus is classified as a metabolic disease, while hypertension and coronary artery disease are considered cardiovascular diseases. With the KBs store the patient’s electronic health record (EHR) and medical abbreviations like “ED” for emergency department, the extracted message serves as a medically meaningful and compressed representation of the patient’s historical conditions, reducing both data volume and transmission delay. The integration of fluid antennas enhances semantic communication by mitigating semantic errors and ensuring accurate message interpretation. Additionally, computing plays a vital role in processing sensed data, simulating physical behaviours, and extracting high-level semantic features necessary for efficient and meaningful communication.
@@
Capital boldface letters (e.g., $\mathbf{A}$) denote matrices, while lowercase boldface letters (e.g., $\mathbf{a}$) denote vectors. Scalars are represented using regular lowercase or uppercase fonts. The sets $\mathbb{C}$, $\mathbb{C}^{n\times 1}$, and $\mathbb{C}^{m\times n}$ represent a complex number, a complex vector of length $n$, and a complex $m\times n$ matrix, respectively. The identity matrix is denoted by $\mathbf{I}$, and the zero matrix by $\mathbf{0}$. The notations $[\cdot]^{H}$, $\operatorname{Tr}(\cdot)$, $\operatorname{vec}(\cdot)$, and $\text{rank}(\cdot)$ represent the Hermitian transpose, the trace, the vectorisation, and the rank of a matrix, respectively. The $l_{2}$ norm is denoted by |||||$\cdot$||| | ⋅ | |. The symbol $\succeq$ indicates positive semi-definiteness, and $\otimes$ denotes the Kronecker product. Finally, $\mathcal{CN}(0,\sigma^{2})$ represents a standard complex Gaussian distribution with zero mean and variance $\sigma^{2}$.
@@ II System Model
with $[\mathbf{x}_{t},\;\mathbf{z}_{t}]$ collects the non-repeating $x$ and $z$ coordinates of all FAs. Similarly, for the coordinates of all FPAs, let $\mathbf{x}_{r}$ and $\mathbf{z}_{r}$ respectively collect the $x$ and $z$ coordinates. Then, we define the collection of FPA coordinates $\mathbf{v}\in\mathbb{C}^{1\times\left(2\times N_{\text{rx}}\times N_{\text{rz}}\right)}$ as
@@
and the non-repeating $x$ and $z$ coordinates of all FPAs are collected by $[\mathbf{x}_{r},\;\mathbf{z}_{r}]$.
The BS communicates with $K$ communication users (CUs), and each CU $k\in\mathcal{K}$ is equipped with a single antenna. To ensure the feasibility of beamforming design in a multi-user communication scenario, we set $K\leq(N_{\text{tx}}\times N_{\text{tz}})$. Simultaneously, the BS actively detects $L$ targets, each target $l\in\mathcal{L}$ being considered as an extended target. Alternative sensing models, such as the parametric scatterer model proposed in [38], represent extended targets using a reduced set of key scatterers and offer promising potential for efficient target representation. By capturing the essential scattering structure with fewer parameters, these models enhance computational efficiency without significantly sacrificing accuracy. In this work, we adopt the extended target model as a general framework, rather than focusing on specific parametric instances, to ensure broad applicability in system design. Nevertheless, our approach is inherently flexible and can be readily adapted to parametric models by reducing the number of scatterers per target to match the key scatterers required by a given model.

### v2

@@ I Introduction
With the rapid growth of emerging applications such as virtual/augmented reality (VR/AR) and digital twins (DTs), the integration of wireless sensing and communication is ready to play a pivotal role in enabling next-generation services [olfat2008optimum, shikh2007joint, bobarshad2010low, nehra2010cross, bobarshad2009m, nehra2010spectral, jia2020channel, kobravi2007cross, shadmand2010multi, towhidlou2017improved, shojaeifard2011joint, yang2025toward, yang2025integrated]. Recognizing its transformative potential, the International Telecommunication Union (ITU) has incorporated integrated sensing and communication (ISAC) as one of the six core usage scenarios in its global vision for the sixth-generation (6G) mobile communication systems.
@@
The basic principle of ISAC is to deploy wireless sensing and communication capability through the shared use of wireless resources, including power, frequency bands, beams, and hardware infrastructure. Numerous studies, to list a few, [dong2022sensing, yuan2020bayesian, su2020secure, lu2024random], have demonstrated that ISAC systems can outperform standalone sensing and communication systems in terms of resource efficiency and can unlock synergistic advantages, where sensing enhances communication and communication supports sensing. On the one hand, sensing-assisted communication uses sensing data to provide contextual awareness, which helps optimize communication processes like beamforming, channel estimation, and resource allocation by inferring channel quality and proactively modifying the system’s settings [liu2022integrated]. On the other hand, in a communication-assisted sensing framework, communication devices support sensing by transmitting sensory data and capturing reflected signals, eliminating the need for dedicated sensors. This simplifies network architecture and reduces hardware costs compared to traditional sensor-based networks [zhang2021overview].
Despite significant advancements in ISAC, meeting the exceptionally high data rate requirements and supporting massive connectivity for emerging applications remains a critical challenge. These demands necessitate the deployment of massive antenna arrays or the use of extremely high-frequency bands [liu2023near]. Specifically, when the distance between the transmitter and receiver is less than $\frac{2D^{2}}{\lambda}$ (i.e. the Rayleigh distance), where $D$ represents the array aperture and $\lambda$ is the wavelength, a near-field (NF) model becomes essential to accurately capture the spherical wave propagation characteristics [babu2024symbol]. For instance, with a transmitter aperture of $D=0.5\;\text{m}$ operating at a frequency of $50\;\text{GHz}$, the NF region extends to $83\;\text{m}$.
@@
Several key advantages emerge when considering the NF model. First, unlike far-field (FF) beamforming, which directs beam energy along a fixed path, NF beam focusing utilises spherical wavefronts to concentrate energy at a specific location [cui2022near, zhang2022beam]. This focusing capability enhances the received signal power for intended users. In terms of sensing capabilities, the NF model demonstrates superior performance in target localisation [wang2024cramer]. The spherical wavefront properties enable simultaneous estimation of both angular and range information with high precision, particularly when employing advanced techniques such as multiple signal classification (MUSIC) with fine-grid processing. While Doppler-based methods in FF scenarios can estimate range through velocity measurements, those methods often yield less accurate results, especially in static or low-mobility scenarios where frequency shifts are minimal. Furthermore, the NF beam focusing effect provides enhanced control over the sensing signal distribution, effectively reducing interference in echo signals and improving overall sensing accuracy [zhang2024physical]. This characteristic is particularly valuable in complex environments where multiple line-of-sight (LoS) targets or reflectors may be present. Several studies have explored the NF-ISAC system. For instance, Wang et al. [wang2023near] analysed NF-ISAC systems with multiple communication users and a single target, investigating both fully digital antenna arrays and hybrid digital-analogue configurations. Their findings highlight the performance enhancements enabled by exploiting the additional distance dimension in NF-ISAC compared to FF-ISAC. Zhao et al. [zhao2024modeling] examined the trade-off between communication and sensing functionalities in both downlink and uplink scenarios. Qu et al. [qu2024near] compared NF beam focusing with FF beam steering, demonstrating through simulations that NF beam focusing effectively mitigates co-angle interference while enhancing both communication and sensing performance.
@@
While NF-ISAC offers promising advantages, it faces substantial challenges, particularly in managing the enormous volume of data transmitted and received in real-time scenarios such as autonomous driving. The excessive data volume can lead to increased latency and computational overhead, hindering system performance [gu2023semantic]. Another critical challenge involves potential data privacy risks. As sensing and communication signals are transmitted simultaneously, sensing targets (which are unintended communication users) may inadvertently receive confidential communication messages, raising significant privacy concerns [zhaohui2024secure]. The integration of semantic communication into the NF-ISAC framework presents a promising solution to these challenges. Semantic communication focuses on transmitting the meaning and relevance of information rather than raw data, enabling intelligent prioritisation and compression based on contextual importance [luo2022semantic]. For example, in autonomous driving, instead of transmitting comprehensive sensing results, the system could extract and transmit semantically relevant information, such as the trajectories of critical objects or potential hazards, to other roadside units. This semantic-aware approach significantly reduces the data burden and enhances system efficiency by ensuring that limited resources are allocated to transmitting the most meaningful and actionable information. To enable semantic communication, both transmitters and receivers must establish shared knowledge bases (KBs) containing essential information accessible to both parties [xu2023edge, gunduz2022beyond]. Crucially, the security and integrity of these KBs are central to maintaining the privacy and security of semantic communication. Regulations such as the General Data Protection Regulation (GDPR), which governs the handling of personal data within the EU and EEA, provide a legal and procedural framework for protecting sensitive data [do2025security, agarwal2022gdpr]. By adhering to GDPR principles or similar regulatory frameworks, KBs can be safeguarded against unauthorised access and cyber-attacks [guo2024survey, won2024resource]. As a result, semantic communication inherently enhances data privacy: since only high-level semantic representations are transmitted, and successful decoding requires access to valid KBs, unintended users are effectively excluded from recovering the original message. Moreover, semantic communication facilitates efficient multi-modal data transmission, meeting the diverse requirements of different communication users [zhang2024unified]. Despite these advantages of semantic communication, existing research on semantic communication, such as [lu2022rethinking, xie2022task, xie2021deep], has largely overlooked its integration into the NF-ISAC framework. This research gap represents a significant opportunity for advancing NF-ISAC systems by leveraging the benefits of semantic communication to address data overload, privacy concerns, and multi-modal data requirements.
@@
Nonetheless, integrating semantic communication into the NF-ISAC framework introduces specific challenges, particularly due to the interference caused by the sensing signals. Unlike classical communication, which focuses on bit-level accuracy, semantic communication aims to preserve the intended meaning or context of the transmitted information. This makes it inherently more sensitive to channel conditions and interferences. Under poor channel conditions, not only can the data become corrupted, but the semantic meaning of a message may also be distorted—a phenomenon referred to as semantic noise [luo2022semantic]. Semantic noise occurs when the receiver misinterprets the message, deviating from the sender’s original intent, which can lead to misunderstandings or incorrect decisions. For example, suppose a sender intends to transmit the message “Alice’s new bike was broken”. In a traditional communication system, a noisy channel might corrupt it to “Alace’s new baki was brakon”, maintaining the structure but affecting intelligibility. In contrast, a semantic communication system might deliver “Alisa’s bike is broken,” preserving a plausible meaning but deviating from the exact intended context. Therefore, developing strategies to mitigate semantic noise is essential for maintaining reliability in semantic communication systems.
@@
One promising technique to stabilise the channel condition is to use the fluid antennas (FAs), which are flexible in terms of their positions to reconfigure the radiation characteristics [ghadi2024physical, zhou2024near, zhu2023movable]. Compared to fixed-position antennas (FPAs), the FAs have a greater degree of freedom (DoF) to explore the channel variations to adapt to the changing channel conditions. Various studies have highlighted the significant advantages of FAs in wireless communication systems. For instance, Zhu et al. [zhu2023modeling] compared the communication performance of FPAs and FAs by analysing the maximum channel gain achieved in deterministic and stochastic channel environments. Their findings show that FAs outperform FPAs, especially in scenarios with an increasing number of channel paths, due to their ability to exploit more pronounced small-scale fading effects in the spatial domain. Similarly, Wang et al. [wang2024fluid] explored the role of FAs in ISAC systems within a multi-user multiple-input multiple-output (MU-MIMO) downlink framework. In this study, a base station (BS) equipped with FAs demonstrated enhanced beamforming precision, reduced inter-user interference, and improved overall system capacity. These advancements illustrate the potential of FAs to optimise both communication and sensing performance in complex multi-user scenarios. Despite the substantial progress made in demonstrating the benefits of FAs for improving communication performance in conventional systems, their applications in semantic communication and NF-ISAC systems have received minimal attention.
@@
In [yang2024secure], the framework of integrated sensing, computing, and semantic communication (ISCSC) was proposed. However, the impact of NF effects and the potential advantages of incorporating FAs into the system were not explored. This limitation highlights a significant research gap, motivating the development of fluid antenna-enabled NF-ISCSC (FA-enabled NF-ISCSC) systems. Compared to the ISAC design, the proposed framework achieves higher data rates without sacrificing sensing performance, or alternatively, enhances sensing accuracy while maintaining the same data rate. These advantages make it especially suitable for applications such as vehicular networks and digital twins. In vehicular networks, particularly within the intelligent transportation system (ITS), sensing enables roadside units (RSUs) and vehicles to acquire essential environmental information, such as traffic signs, road conditions, and dynamic obstacles [yang2022semantic]. Semantic communication ensures that only task-relevant and context-specific information, such as the sudden appearance of a pedestrian, is exchanged among RSUs and between RSUs and vehicles. For example, rather than transmitting raw sensor data, an RSU can convey a high-level semantic message like “vehicle merging from left at 25 m/s”, thereby reducing the communication load and improving the accuracy and timeliness of information delivery. The use of FAs further enhances communication reliability and adaptability by reducing semantic errors. Computing plays a critical role by processing sensed data for real-time scene understanding, including object detection, tracking, and trajectory prediction [wen2024survey]. Additionally, computing enables semantic communication by extracting and prioritising key features from sensory data for efficient and meaningful transmission. In the context of digital twins, sensing data collected by end devices enables real-time perception of the physical entity’s state. In the uplink scenario, semantic communication allows the end device to interpret and filter out irrelevant information prior to the transmission, thereby reducing bandwidth consumption and minimising latency when uploading to the digital twin server [jagatheesaperumal2023semantic]. In the downlink, the server can generate and transmit semantic information based on environmental context, thus alleviating downlink resource congestion. For example, in the context of a digital twin for eHealth, instead of transmitting the full message “The patient is a 68-year-old male with a medical history of Type 2 diabetes mellitus, hypertension, and coronary artery disease. He presented to the emergency department with chest pain that radiated to his left arm”, the end device may send a high-level semantic interpretation such as “68-year-old male with chronic cardiovascular and metabolic diseases presented with radiating chest pain to the ED”. In this semantically extracted message, the word “chronic” captures the notion of a long-standing medical history, Type 2 diabetes mellitus is classified as a metabolic disease, while hypertension and coronary artery disease are considered cardiovascular diseases. With the KBs store the patient’s electronic health record (EHR) and medical abbreviations like “ED” for emergency department, the extracted message serves as a medically meaningful and compressed representation of the patient’s historical conditions, reducing both data volume and transmission delay. The integration of fluid antennas enhances semantic communication by mitigating semantic errors and ensuring accurate message interpretation. Additionally, computing plays a vital role in processing sensed data, simulating physical behaviours, and extracting high-level semantic features necessary for efficient and meaningful communication.
@@
Capital boldface letters (e.g., $\mathbf{A}$) denote matrices, while lowercase boldface letters (e.g., $\mathbf{a}$) denote vectors. Scalars are represented using regular lowercase or uppercase fonts. The sets $\mathbb{C}$, $\mathbb{C}^{n\times 1}$, and $\mathbb{C}^{m\times n}$ represent a complex number, a complex vector of length $n$, and a complex $m\times n$ matrix, respectively. The identity matrix is denoted by $\mathbf{I}$, and the zero matrix by $\mathbf{0}$. The notations $\cdot]^{H}$, $\operatorname{Tr}(\cdot)$, $\operatorname{vec}(\cdot)$, and $\text{rank}(\cdot)$ represent the Hermitian transpose, the trace, the vectorisation, and the rank of a matrix, respectively. The $l_{2}$ norm is denoted by $||\cdot||$. The symbol $\succeq$ indicates positive semi-definiteness, and $\otimes$ denotes the Kronecker product. Finally, $\mathcal{CN}(0,\sigma^{2})$ represents a standard complex Gaussian distribution with zero mean and variance $\sigma^{2}$.
@@ II System Model
with $\mathbf{x}_{t},\;\mathbf{z}_{t}]$ collects the non-repeating $x$ and $z$ coordinates of all FAs. Similarly, for the coordinates of all FPAs, let $\mathbf{x}_{r}$ and $\mathbf{z}_{r}$ respectively collect the $x$ and $z$ coordinates. Then, we define the collection of FPA coordinates $\mathbf{v}\in\mathbb{C}^{1\times\left(2\times N_{\text{rx}}\times N_{\text{rz}}\right)}$ as
@@
and the non-repeating $x$ and $z$ coordinates of all FPAs are collected by $\mathbf{x}_{r},\;\mathbf{z}_{r}]$.
The BS communicates with $K$ communication users (CUs), and each CU $k\in\mathcal{K}$ is equipped with a single antenna. To ensure the feasibility of beamforming design in a multi-user communication scenario, we set $K\leq(N_{\text{tx}}\times N_{\text{tz}})$. Simultaneously, the BS actively detects $L$ targets, each target $l\in\mathcal{L}$ being considered as an extended target. Alternative sensing models, such as the parametric scatterer model proposed in [liu2024crb], represent extended targets using a reduced set of key scatterers and offer promising potential for efficient target representation. By capturing the essential scattering structure with fewer parameters, these models enhance computational efficiency without significantly sacrificing accuracy. In this work, we adopt the extended target model as a general framework, rather than focusing on specific parametric instances, to ensure broad applicability in system design. Nevertheless, our approach is inherently flexible and can be readily adapted to parametric models by reducing the number of scatterers per target to match the key scatterers required by a given model.